OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (5, 7) and (8, 22)?

OpenStudy (hoblos):

first, lets calculate the slope between the two points do you know how to calculate the slope ?

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (hoblos):

correct so what would it be for the given points ?

OpenStudy (anonymous):

I got 17/3

OpenStudy (hoblos):

22-7 = ?

OpenStudy (anonymous):

I mean 15/3

OpenStudy (anonymous):

simplified to 5

OpenStudy (hoblos):

correct, now we have to apply the formula y - y1 = m( x - x1) to get the equation in point slope form, and then we'll transform it to the slope-intercept form

OpenStudy (hoblos):

can you apply y - y1 = m( x - x1) please

OpenStudy (anonymous):

y-7=5x-25 y=5-18

OpenStudy (hoblos):



OpenStudy (anonymous):

yes, sorry

OpenStudy (anonymous):

lol

OpenStudy (hoblos):

then the answer is correct good work :)

OpenStudy (anonymous):

thank you :)

OpenStudy (hoblos):

any time :)